### v1

# GASE: Generatively Augmented Sentence Encoding

Abstract
We propose an approach to enhance sentence embeddings by applying generative text models for data augmentation at inference time. Unlike conventional data augmentation that utilises synthetic training data, our approach does not require access to model parameters or the computational resources typically required for fine-tuning state-of-the-art models. Generatively Augmented Sentence Encoding uses diverse linguistic synthetic variants of input texts generated by paraphrasing, summarising, or extracting keywords, followed by pooling the original and synthetic embeddings. Experimental results on the Massive Text Embedding Benchmark for Semantic Textual Similarity (STS) demonstrate performance improvements across a range of embedding models using different generative models for augmentation. We find that generative augmentation leads to larger performance improvements for embedding models with lower baseline performance. These findings suggest that integrating generative augmentation at inference time adds semantic diversity and can enhance the robustness and generalizability of sentence embeddings for embedding models. Our results show that the degree to which generative augmentation can improve STS performance depends not only on the embedding model but also on the dataset. From a broader perspective, the approach allows trading training for inference compute.
GASE: Generatively Augmented Sentence Encoding
Manuel Frank mfr-research@protonmail.com                        Haithem Afli Department of Computer Science Munster Technological University Haithem.Afli@mtu.ie
Figure 1: Approach for Generatively Augmented Sentence Encoding.
1 Introduction
Representation learning has emerged as a fundamental technique in natural language processing (NLP). Still, the quality and robustness of embeddings largely depend on the richness and diversity of the training data. To derive sentence embeddings from BERT (Devlin et al., 2019) word embeddings, these can be averaged or the embedding of the CLS token [^1] CLS tokens are special tokens placed at the beginning of each input example in a BERT model, providing a representation of the entire input for use in classification tasks (Devlin et al., 2019). being used. Based on encoder-style transformers, such as BERT, Sentence Transformers (STs) (Reimers and Gurevych, 2020) were trained to directly output sentence embeddings by calculating an attention-weighted average of the word embeddings. These sentence embeddings can be used in pairwise tasks, e.g. paraphrase identification, or non-pairwise tasks, e.g. sentiment analysis. Recent advancements in generative models, such as OpenAI’s GPT-4 (OpenAI, 2023), Anthropic’s Claude 3 models (Anthropic, 2024) and Google’s Gemini models (Gemini Team et al., 2024), have shown remarkable capabilities in generating human-like text. Moreover, generative models like ChatGPT have been demonstrated to be capable of paraphrasing, e.g. for data augmentation Dai et al. (2023). The authors use ChatGPT to rephrase a given text to generate synthetic data which is used to train a downstream BERT model generating embeddings for text classification. To augment generative models, the integration with representation learning has been explored, e.g. in Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020). In contrast, we introduce Generatively Augmented Sentence Encoding (GASE), an approach to augment embedding models using generative Large Language Models (LLMs). Instead of generating synthetic data for pre-training or fine-tuning, we propose an augmentation approach which is applied at inference time by generating one or more textual variants of input data and pooling the embeddings of the original input text and the synthetic texts. The underlying hypothesis of our work on generative augmentation is that generative models can introduce semantic diversity which is not modelled by current embedding models without generative augmentation and, therefore, can benefit the performance of downstream tasks using the embeddings. Our method involves three generative tasks to generate the synthetic data from the original data: paraphrasing, summarising, and extracting keywords. To the best of our knowledge, this is the first work proposing generative models to augment sentence encoders at inference time.
2 Approach
As illustrated in Figure 1, GASE consists of the following steps:
1. Generative Augmentation. In this step $k$ generative text models are used to produce $k$ variations of the original input sequence. We propose three augmentation functions: Paraphrasing, which provides a semantically equivalent but lexically or syntactically different variation of the input text. Summarising which produces a shorter output text that captures the most important information of the input text. And Extracting Keywords which lists the most relevant words from a given text. An example for each is provided in Figure 2.
Figure 2: Augmentation examples for paraphrasing, summarising, and extracting keywords.
We evaluate $k\in\{0,1,2\}$ using the generative LLMs GPT-3.5 Turbo (OpenAI, 2024b), Reka-Flash (Reka, 2024), and GPT-4o mini  (OpenAI, 2024a) with temperature and top_p set to 00 and $1$ respectively. As a baseline, we use $k=0$, i.e., no augmentation. The prompts can be found in Appendix A. Model-specific post-processing is performed to remove meta-text that is not used for augmentation.
2. Sentence Embeddings. For the original sequence and each of the $k$ synthetic variations an embedding is inferred. We evaluated the following embedding models which include non-transformer baselines, locally run ST models, and SOTA models accessed via an API using default hyperparameter values: glove.840B.300d (Pennington et al., 2014), bert-large-cased (Devlin et al., 2019), all-MiniLM-L6-v2 Sentence Transformers (2024), all-mpnet-base-v2 Sentence Transformers (2024), embed-english-light-v3.0 Cohere (2024), embed-english-v3.0 Cohere (2024), voyage-2 (Voyage AI, 2024), voyage-lite-02-instruct,  (Voyage AI, 2024), voyage-large-2 (Voyage AI, 2024), voyage-large-2-instruct (Voyage AI, 2024), mxbai-embed-large-v1 Mixedbread (2024).
3. Pooling. The embeddings of the original text and the $k$ augmented texts are pooled by computing their arithmetic mean.
We evaluate our approach on the English subtasks of the MTEB STS task [^2] This includes STSBenchmark (Cer et al., 2017a), STS12 (Agirre et al., 2012), STS13 (Agirre et al., 2013), STS14 Agirre et al. (2014), STS15 (Agirre et al., 2015), STS16 (Agirre et al., 2016), STS17 (Cer et al., 2017b), STS22 (Chen et al., 2022), SICK-R Marelli et al. (2014), and BIOSSES (Soğancıoğlu et al., 2017). According to the original paper (Muennighoff et al., 2023), MTEB also includes the STS11 dataset. However, the MTEB leaderboard (Hugging Face, 2024) does not. To make our results comparable to the leaderboard we excluded STS11. using cosine similarity and Spearman’s rank correlation [^3] See  Reimers et al. (2016) for why Spearman’s rank correlation is preferable over Pearson correlation for STS tasks..
3 Results
Table 1 shows the results for different augmentation strategies using GPT-3.5 Turbo vs. the respective baseline. (See Table 4 in the Appendix for the results per dataset.) In terms of average scores across datasets, all embedding models benefit from some form of augmentation.
Embedding
model
Augmentation
strategy
Avg.    STS22
glove. 840B.300d - 57.86    54.08
paraphrasing 59.98    57.78
summarising 60.64    61.21
extracting keywords 60.82    56.94
bert- large-cased - 60.78    55.48
paraphrasing 66.29    58.57
summarising 66.89    63.47
extracting keywords 62.62    54.64
all-MiniLM- L6-v2 - 78.91    67.26
paraphrasing 80.32    68.47
summarising 79.90    69.27
extracting keywords 77.96    64.57
all-mpnet- base-v2 - 80.28    68.00
paraphrasing 81.39    69.01
summarising 81.48    70.12
extracting keywords 79.95    66.19
embed-english- light-v3.0 - 78.75    67.88
paraphrasing 80.30    68.42
summarising 80.21    68.97
extracting keywords 78.53    67.27
embed- english-v3.0 - 81.25    68.22
paraphrasing 82.10    68.63
summarising 82.08    70.29
extracting keywords 81.26    66.91
voyage-2 - 82.50    65.26
paraphrasing 83.38    66.80
summarising 82.74    68.50
extracting keywords 81.21    65.29
voyage-lite- 02-instruct - 85.95    78.62
paraphrasing 85.99    79.08
summarising 85.43    79.87
extracting keywords 83.91    76.60
voyage- large-2 - 83.60    63.97
paraphrasing 83.92    64.69
summarising 83.66    65.90
extracting keywords 81.98    63.68
voyage-large- 2-instruct - 84.61    66.59
paraphrasing 84.80    67.72
summarising 84.54    68.66
extracting keywords 83.40    65.73
mxbai-embed- large-v1 - 84.63    68.73
paraphrasing 85.01    70.03
summarising 84.45    70.36
extracting keywords 83.42    67.83
Table 1: Average scores on the MTEB datasets and STS22 individually for all embedding models with and without generative augmentation using GPT-3.5 Turbo (Spearman’s rank correlation in %, bold: highest scores).
The largest performance increases are observed for BERT (+$6.11$pp) and GloVe (+$2.96$pp). GloVe is also the only embedding model achieving its best results using keyword extraction. Moreover, BERT and MPNET STs benefit most from summarisation. In terms of datasets, all embedding models scored lowest on STS22.
Dataset Original
Para-
phrasing
Summa-
rising
Extracting
keywords
STSB 10.1 11.5 10.7 5.7
STS12 11.1 11.6 11.1 6.1
STS13 9.0 10.5 11.1 5.4
STS14 9.3 11.0 11.4 5.7
STS15 10.6 12.0 11.5 5.7
STS16 11.6 12.6 12.8 5.7
STS17 8.7 9.6 9.0 4.5
STS22 477.2 216.5 58.4 103.8
SICK-R 9.6 10.0 9.3 4.7
BIOSSES 24.5 25.2 19.2 13.9
Average 58.2 33.1 16.4 16.1
Table 2: Mean word count (using GPT-3.5 Turbo for generative augmentation).
As expected, summarising and extracting keywords significantly reduced the mean word count on average across datasets (see Table 2). While paraphrases for all datasets except STS22 are close to the word count of the original texts, GPT-3.5 Turbo does not maintain STS22’s high word count.
Since paraphrase augmentation provided the largest improvements for most models, we evaluated it for all generative models (see Table 3 and for results per dataset Table 5 in the Appendix). Using a single generative model, GPT-3.5 Turbo achieved the best results for all embedding models except GloVe and Voyage-Large-2 which performed better using Reka-Flash. Hence, we ran experiments with these two models generating $k=2$ paraphrases.
Across embedding models, the use of these two generative models led to higher scores than a single generative model. The largest performance gains were achieved with BERT (+$7.23$pp) and GloVe (+$4.11$pp). MiniLM ST, MPNet ST, Cohere’s Embed models and Voyage-2 demonstrated maximum improvements between $1.28$ and $1.79$pp while the remaining embedding models showed performance gains below $1$pp. Notably, the results with $k=2$ surpass the mean between the scores of employing GPT-3.5 Turbo and Reka Flash individually.
Embedding
model
Generative model
Avgerage
score
glove. 840B.300d - 57.86
gpt-3.5-turbo 59.98
reka-flash 60.73
GPT-4o mini 60.04
gpt-3.5-turbo, reka-flash 61.97
bert- large-cased - 60.78
gpt-3.5-turbo 66.29
reka-flash 65.76
GPT-4o mini 65.01
gpt-3.5-turbo, reka-flash 68.01
all-MiniLM- L6-v2 - 78.91
gpt-3.5-turbo 80.32
reka-flash 79.74
GPT-4o mini 79.81
gpt-3.5-turbo, reka-flash 80.64
all-mpnet- base-v2 - 80.28
gpt-3.5-turbo 81.39
reka-flash 81.01
GPT-4o mini 81.07
gpt-3.5-turbo, reka-flash 81.72
embed-english- light-v3.0 - 78.75
gpt-3.5-turbo 80.30
reka-flash 79.72
GPT-4o mini 79.90
gpt-3.5-turbo, reka-flash 80.54
embed- english-v3.0 - 81.25
gpt-3.5-turbo 82.10
reka-flash 82.00
GPT-4o mini 81.82
gpt-3.5-turbo, reka-flash 82.53
voyage-2 - 82.50
gpt-3.5-turbo 83.38
reka-flash 83.37
GPT-4o mini 83.07
gpt-3.5-turbo, reka-flash 83.80
voyage-lite- 02-instruct - 85.95
gpt-3.5-turbo 85.99
reka-flash 85.92
GPT-4o mini 85.97
gpt-3.5-turbo, reka-flash 86.05
voyage- large-2 - 83.60
gpt-3.5-turbo 83.92
reka-flash 83.98
GPT-4o mini 83.95
gpt-3.5-turbo, reka-flash 84.12
voyage-large- 2-instruct - 84.61
gpt-3.5-turbo 84.80
reka-flash 84.68
GPT-4o mini 84.78
gpt-3.5-turbo, reka-flash 84.86
mxbai-embed- large-v1 - 84.63
gpt-3.5-turbo 85.01
reka-flash 84.80
GPT-4o mini 84.90
gpt-3.5-turbo, reka-flash 85.17
Table 3: Average scores on the MTEB datasets with and without paraphrase augmentation using GPT-3.5 Turbo, GPT-4o mini, and Reka-Flash (Spearman’s rank correlation in %, bold: highest scores).
Figure 3: Score without augmentation vs. max. score with augmentation per embedding model.
4 Discussion
Overall, lower non-augmented performance was associated with larger performance improvements through generative augmentation as visualized in Figure 3. We believe this relationship holds because weaker embedding models are inherently less capable of modelling diverse semantics and, therefore, benefit more from the semantic diversity introduced by generative models. Moreover, embedding models with lower baseline performance also gained more from the ensemble effect of using $k=2$ models for augmentation compared to a single generative model. The effectiveness of keyword extraction for GloVe may be explained by the model’s limited capability to handle complex semantics as its sentence embeddings are derived by calculating an unweighted average of word embeddings. Therefore, the reduction to keywords might help the model to reduce noise.
We believe the lower scores on STS22 can be explained by its higher word count. Unexpectedly, generating summaries did not significantly reduce the mean word count on datasets other than STS22 (and BIOSSES, to some extent). Since most MTEB STS datasets include shorter sentences, sentence encoders might be optimized for shorter sentence lengths. Our results demonstrate that the best augmentation strategy for a given embedding model is not identical for all datasets.
5 Conclusion
We propose Generatively Augmented Sentence Encoding (GASE), a novel approach that augments sentence embeddings by paraphrasing, summarising, or extracting keywords with generative models at inference time. On the MTEB STS task, GASE leads to substantial improvements for embedding models with lower baseline performance and incremental improvements for SOTA embedding models. GASE is widely applicable as it only requires access to the outputs of the underlying embedding and generative models.
6 Limitations
This study presents several limitations:
  1. 1.
    While deterministic embedding models such as GloVe yield consistent results, other models like GPT-3.5 Turbo examined in this research exhibit stochastic behaviour. Consequently, additional experiments are necessary to corroborate our findings, as LLMs can produce diverse outcomes, thereby limiting the conclusiveness of a single experimental run (Reimers and Gurevych, 2018). This caveat is particularly pertinent to observations based on marginal performance differentials.
  2. 2.
    Our investigation was confined to the evaluation of GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini for generative augmentation. Alternative LLMs, including the Claude and Gemini model families, may potentially offer more substantial improvements.
  3. 3.
    Furthermore, our experimental design was restricted to $k\in\{0,1,2\}$. While values of $k>2$ were not explored, we posit that such configurations may be impractical for real-world applications due to prohibitive computational costs.
  4. 4.
    Summaries typically comprise $50\%$ or fewer words relative to its source text (Radev et al., 2002). However, due to the concise nature of the texts within the examined STS datasets the summaries generated in this work approximately maintain the original word count. (Except for STS22 which contains longer text sequences.) Hence, the efficacy of augmentation through summarisation may be more pronounced when applied to datasets comprising longer textual inputs (e.g., paragraphs) compared to the predominantly short text sequences examined within the scope of this study.
  5. 5.
    Our experiments are limited to the English language and the effectiveness of GASE for other languages requires further investigation. Since embedding models with weaker baseline performance showed larger improvements through generative augmentation, we hypothesise that GASE can provide larger performance improvements for languages for which the performance of embedding models is below English.
Acknowledgments
Code development for this work has been assisted by GPT-3.5 Turbo, GPT-4 Omni, Claude 3.5 Sonnet, Gemini Pro 1.5, and Github Copilot.
Appendix A Prompts
Paraphrasing with GPT-3.5 Turbo and GPT-4o mini :
"Rephrase the following text while maintaining its original meaning. If the text contains only a single word, provide a definition or a synomym. When done, check and make sure that the length of the original is approximately maintained. Text:"
Paraphrasing with Reka-Flash:
"Rephrase the following text while maintaining its original meaning. Do not provide multiple alternatives. Before you reply, remove any explanations. Do only reply with the paraphrased text. If the text contains only a single word, provide a definition or a synomym. Text:"
summarising with GPT-3.5 Turbo:
"summarise the following text. Do not include any meta text and only output the summary. If the text is too short to summarise, paraphrase it instead. Text:"
Extracting keywords with GPT-3.5 Turbo:
"Extract the keywords from the following sentence. Do NOT inlcude any commas or fullstops in your response and do not start your answer with "keywords". Text: "
Appendix B Additional Experimental Results
Embedding model Augmentation Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
glove.840B.300d - 57.86 50.73 57.50 70.98 60.69 70.85 63.85 62.05 54.08 55.42 32.45
paraphrasing 59.98 56.78 57.33 71.17 61.16 71.52 67.66 62.28 57.78 58.92 35.21
summarising 60.64 57.67 55.00 70.24 61.18 71.64 66.81 67.08 61.21 60.92 34.65
extracting keywords 60.82 59.26 58.09 71.12 60.34 72.44 65.31 75.92 56.94 55.34 33.50
bert-large-cased - 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
paraphrasing 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.79 64.56
summarising 66.89 68.23 51.39 70.58 61.98 74.27 70.60 74.48 63.47 70.74 63.23
extracting keywords 62.62 63.03 53.50 63.32 58.51 72.62 68.13 71.99 54.64 58.24 62.22
all-MiniLM-L6-v2 - 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
paraphrasing 80.32 82.33 73.33 83.33 77.85 86.23 79.88 87.02 68.47 79.63 85.07
summarising 79.90 82.17 72.14 82.89 77.24 86.34 78.99 87.70 69.27 79.86 82.41
extracting keywords 77.96 81.99 73.29 81.14 75.77 85.87 78.70 87.45 64.57 70.92 79.88
all-mpnet-base-v2 - 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
paraphrasing 81.39 84.10 74.04 85.52 80.52 86.82 81.15 89.08 69.01 81.25 82.39
summarising 81.48 83.94 73.76 85.54 79.78 86.73 80.20 90.41 70.12 81.32 82.96
extracting keywords 79.95 82.56 74.82 84.82 78.74 86.31 79.01 90.23 66.19 77.07 79.74
embed-english-light-v3.0 - 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
paraphrasing 80.30 84.65 74.65 80.73 79.41 85.45 81.14 87.45 68.42 78.75 82.36
summarising 80.21 84.08 73.45 81.36 78.41 86.26 81.52 88.19 68.97 79.08 80.79
extracting keywords 78.53 82.81 75.03 80.17 78.29 84.14 76.38 88.14 67.27 74.45 78.64
embed-english-v3.0 - 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
paraphrasing 82.10 86.69 76.37 83.71 81.29 87.96 83.95 88.47 68.63 78.21 85.75
summarising 82.08 86.47 74.40 84.41 80.76 88.49 83.55 89.08 70.29 78.06 85.25
extracting keywords 81.26 85.50 74.77 84.66 79.98 87.72 81.09 88.75 66.91 77.53 85.68
voyage-2 - 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
paraphrasing 83.38 87.17 79.23 87.13 81.67 88.57 86.27 89.07 66.80 80.33 87.56
summarising 82.74 86.51 76.81 86.25 80.89 88.46 84.88 89.10 68.50 80.48 85.55
extracting keywords 81.21 85.71 77.76 86.17 79.63 87.49 84.37 88.66 65.29 72.06 85.03
voyage-lite-02-instruct - 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
paraphrasing 85.99 88.81 84.47 89.05 86.68 89.52 86.53 86.79 79.08 79.09 89.89
summarising 85.43 87.93 82.77 88.06 85.05 88.98 85.49 86.70 79.87 79.16 90.32
extracting keywords 83.91 86.75 83.82 87.97 84.47 88.98 84.55 86.28 76.60 70.77 88.88
voyage-large-2 - 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
paraphrasing 83.92 87.62 79.18 87.66 83.60 88.91 86.00 90.22 64.69 81.00 90.29
summarising 83.66 87.56 77.83 86.93 82.88 88.93 84.88 90.49 65.90 81.09 90.15
extracting keywords 81.98 85.92 78.00 86.45 81.33 88.28 84.44 89.90 63.68 72.50 89.33
voyage-large-2-instruct - 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
paraphrasing 84.80 89.31 77.61 88.40 86.62 91.17 86.20 89.58 67.72 82.83 88.58
summarising 84.54 88.84 76.27 88.29 85.56 91.06 85.24 89.40 68.66 83.28 88.78
extracting keywords 83.40 87.38 76.63 88.58 85.55 90.51 84.03 88.57 65.73 76.58 90.46
mxbai-embed-large-v1 - 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
paraphrasing 85.01 89.09 80.83 89.43 85.88 89.46 86.94 88.75 70.03 82.67 87.02
summarising 84.45 88.60 78.51 89.07 84.82 89.09 85.96 88.43 70.36 82.82 86.82
extracting keywords 83.42 87.43 78.64 88.75 83.82 88.44 84.84 87.40 67.83 78.17 88.90
Table 4: Scores per dataset with and without generative augmentation using GPT-3.5 Turbo (Spearman’s rank correlation in %, bold: highest scores).
Embedding model Gen model Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
glove.840B.300d - 57.86 50.73 57.50 70.98 60.69 70.85 63.85 62.05 54.08 55.42 32.45
gpt-3.5-turbo 59.98 56.78 57.33 71.17 61.16 71.52 67.66 62.28 57.78 58.92 35.21
reka-flash 60.73 55.02 59.53 71.09 61.22 72.81 65.20 64.27 60.11 61.65 36.43
GPT-4o mini 60.04 55.09 60.12 72.65 62.45 71.60 66.49 63.39 55.48 60.36 32.76
gpt-3.5-turbo, reka-flash 61.97 59.46 60.06 72.01 62.68 72.46 67.62 65.43 61.46 61.51 36.95
bert-large-cased - 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
gpt-3.5-turbo 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.79 64.56
reka-flash 65.76 65.20 52.92 68.55 59.77 72.67 71.80 71.46 58.93 70.09 66.19
GPT-4o mini 65.01 65.37 54.27 68.60 58.18 72.26 71.12 73.54 55.72 70.50 60.54
gpt-3.5-turbo, reka-flash 68.01 69.39 56.86 70.04 62.30 74.61 73.86 74.52 60.49 71.32 66.72
all-MiniLM-L6-v2 - 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
gpt-3.5-turbo 80.32 82.33 73.33 83.33 77.85 86.23 79.88 87.02 68.47 79.63 85.07
reka-flash 79.74 81.30 73.52 82.42 76.83 85.72 79.02 87.01 68.83 79.66 83.11
GPT-4o mini 79.81 81.97 74.03 83.31 76.70 86.03 79.19 87.69 67.01 80.08 82.11
gpt-3.5-turbo, reka-flash 80.64 82.30 74.59 83.72 78.11 86.20 79.95 87.23 69.23 80.12 84.97
all-mpnet-base-v2 - 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
gpt-3.5-turbo 81.39 84.10 74.04 85.52 80.52 86.82 81.15 89.08 69.01 81.25 82.39
reka-flash 81.01 83.72 74.27 85.12 79.33 86.41 80.41 90.17 68.42 81.37 80.87
GPT-4o mini 81.07 83.75 74.19 85.19 79.28 86.45 80.47 90.89 67.81 81.23 81.43
gpt-3.5-turbo, reka-flash 81.72 84.43 75.46 86.12 80.63 87.00 81.21 89.37 69.33 81.55 82.13
embed-english-light-v3.0 - 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
gpt-3.5-turbo 80.30 84.65 74.65 80.73 79.41 85.45 81.14 87.45 68.42 78.75 82.36
reka-flash 79.72 83.51 75.74 78.60 78.38 84.63 80.09 87.98 69.06 78.76 80.42
GPT-4o mini 79.90 84.14 75.76 79.65 77.89 85.12 79.83 87.90 67.91 78.68 82.10
gpt-3.5-turbo, reka-flash 80.54 84.48 76.45 80.61 79.65 85.54 81.21 87.60 69.23 79.06 81.55
embed-english-v3.0 - 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
gpt-3.5-turbo 82.10 86.69 76.37 83.71 81.29 87.96 83.95 88.47 68.63 78.21 85.75
reka-flash 82.00 86.31 76.91 82.22 80.48 87.82 83.69 88.93 69.70 78.06 85.90
GPT-4o mini 81.82 86.52 76.80 82.57 80.11 87.77 83.72 89.42 68.36 77.90 85.05
gpt-3.5-turbo, reka-flash 82.53 86.68 77.88 83.56 81.64 88.05 84.13 88.59 69.87 78.44 86.48
voyage-2 - 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
gpt-3.5-turbo 83.38 87.17 79.23 87.13 81.67 88.57 86.27 89.07 66.80 80.33 87.56
reka-flash 83.37 86.93 79.65 87.83 81.79 88.48 86.18 88.98 67.42 80.67 85.72
GPT-4o mini 83.07 86.89 79.59 87.27 81.23 88.46 85.81 89.37 65.25 80.85 86.02
gpt-3.5-turbo, reka-flash 83.80 87.27 80.62 87.99 82.40 88.51 86.40 88.77 67.94 80.87 87.28
voyage-lite-02-instruct - 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
gpt-3.5-turbo 85.99 88.81 84.47 89.05 86.68 89.52 86.53 86.79 79.08 79.09 89.89
reka-flash 85.92 88.24 84.88 88.79 86.58 89.46 86.28 86.74 78.76 78.97 90.53
GPT-4o mini 85.97 88.25 85.24 89.06 86.62 89.57 85.99 86.90 78.59 79.45 90.08
gpt-3.5-turbo, reka-flash 86.05 88.74 84.83 89.14 86.69 89.36 86.57 86.72 78.81 79.34 90.34
voyage-large-2 - 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
gpt-3.5-turbo 83.92 87.62 79.18 87.66 83.60 88.91 86.00 90.22 64.69 81.00 90.29
reka-flash 83.98 87.32 79.52 88.25 83.49 88.77 85.90 90.40 64.46 81.35 90.37
GPT-4o mini 83.95 87.54 80.02 87.91 83.22 89.11 85.68 90.47 63.99 81.54 90.08
gpt-3.5-turbo, reka-flash 84.12 87.56 80.05 88.31 83.98 88.76 85.97 90.04 64.84 81.55 90.14
voyage-large-2-instruct - 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
gpt-3.5-turbo 84.80 89.31 77.61 88.40 86.62 91.17 86.20 89.58 67.72 82.83 88.58
reka-flash 84.68 88.81 78.41 88.30 86.40 90.95 85.42 89.71 67.04 83.22 88.52
GPT-4o mini 84.78 88.83 78.71 88.43 86.45 91.17 85.89 89.14 67.09 83.19 88.87
gpt-3.5-turbo, reka-flash 84.86 89.16 79.11 88.56 86.67 90.97 85.86 89.67 67.56 83.00 88.08
mxbai-embed-large-v1 - 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
gpt-3.5-turbo 85.01 89.09 80.83 89.43 85.88 89.46 86.94 88.75 70.03 82.67 87.02
reka-flash 84.80 88.93 81.04 89.36 85.51 89.27 86.48 89.04 68.67 82.66 87.07
GPT-4o mini 84.90 88.95 81.08 89.76 85.37 89.43 86.58 89.14 69.10 82.73 86.89
gpt-3.5-turbo, reka-flash 85.17 89.16 81.96 89.37 85.96 89.39 86.79 88.97 69.51 82.72 87.81
Table 5: Scores per dataset with and without paraphrase augmentation using GPT-3.5 Turbo, GPT-4o mini, and Reka-Flash (Spearman’s rank correlation in %, bold: highest scores).