### v2

# GASE: Generatively Augmented Sentence Encoding

Abstract
We propose a training-free approach to improve sentence embeddings leveraging test-time compute by applying generative text models for data augmentation at inference time. Unlike conventional data augmentation that utilises synthetic training data, our approach does not require access to model parameters or the computational resources typically required for fine-tuning state-of-the-art models. Generatively Augmented Sentence Encoding variates the input text by paraphrasing, summarising, or extracting keywords, followed by pooling the original and synthetic embeddings. Experimental results on the Massive Text Embedding Benchmark for Semantic Textual Similarity (STS) demonstrate performance improvements across a range of embedding models using different generative models for augmentation. We find that generative augmentation leads to larger performance improvements for embedding models with lower baseline performance. These findings suggest that integrating generative augmentation at inference time adds semantic diversity and can enhance the robustness and generalisability of sentence embeddings for embedding models. Our results show that performance gains depend on the embedding model and the dataset.
GASE: Generatively Augmented Sentence Encoding
Manuel Frank Department of Computer Science Munster Technological University Manuel.Frank@zohomail.eu          Haithem Afli Department of Computer Science Munster Technological University Haithem.Afli@mtu.ie
Figure 1: Approach for Generatively Augmented Sentence Encoding.
1 Introduction
Figure 2: Augmentation examples for paraphrasing, summarising, and extracting keywords.
Embedding model No augmentation Paraphrasing Summarising
Extracting keywords
Avg. STS22 Avg. STS22 Avg. STS22 Avg. STS22
glove.840B.300d 57.86 54.08 59.98 57.78 60.64 61.21 60.82 56.94
bert-large-cased 60.78 55.48 66.29 58.57 66.89 63.47 62.62 54.64
all-MiniLM-L6-v2 78.91 67.26 80.32 68.47 79.90 69.27 77.96 64.57
all-mpnet-base-v2 80.28 68.00 81.39 69.01 81.48 70.12 79.95 66.19
embed-english-light-v3.0 78.75 67.88 80.30 68.42 80.21 68.97 78.53 67.27
embed-english-v3.0 81.25 68.22 82.10 68.63 82.08 70.29 81.26 66.91
voyage-2 82.50 65.26 83.38 66.80 82.74 68.50 81.21 65.29
voyage-lite-02-instruct 85.95 78.62 85.99 79.08 85.43 79.87 83.91 76.60
voyage-large-2 83.60 63.97 83.92 64.69 83.66 65.90 81.98 63.68
voyage-large-2-instruct 84.61 66.59 84.80 67.72 84.54 68.66 83.40 65.73
mxbai-embed-large-v1 84.63 68.73 85.01 70.03 84.45 70.36 83.42 67.83
llama-3-8b 71.36 33.39 75.28 50.55 76.74 59.44 74.95 51.79
Table 1: Average scores across MTEB STS datasets and STS22 individually for all embedding models with and without generative augmentation using GPT-3.5 Turbo in %, bold: highest average scores for average and STS22 respectively).
Dataset
No Aug-
-mentation
Keyword
Extraction
Random
Keyword
Extraction
STSB 50.73 59.26 23.95
STS12 57.50 58.09 34.78
STS13 70.98 71.12 22.25
STS14 60.69 60.34 28.48
STS15 70.85 72.44 33.86
STS16 63.85 65.31 24.99
STS17 62.05 75.92 36.47
STS22 54.08 56.94 38.44
SICK-R 55.42 55.34 40.04
BIOSSES 32.45 33.50 29.02
Average 57.86 60.82 31.23
Table 2: STS scores per dataset for GloVe embeddings without generative augmentation vs. generative augmentation using LLM-extracted keywords with GPT-3.5 Turbo or random keyword extraction (Spearman’s rank correlation in %, bold: highest scores).
Dataset Original
Para-
phrases
Summa-
ries
Extracted
keywords
STSB 10.1 11.5 10.7 5.7
STS12 11.1 11.6 11.1 6.1
STS13 9.0 10.5 11.1 5.4
STS14 9.3 11.0 11.4 5.7
STS15 10.6 12.0 11.5 5.7
STS16 11.6 12.6 12.8 5.7
STS17 8.7 9.6 9.0 4.5
STS22 477.2 216.5 58.4 103.8
SICK-R 9.6 10.0 9.3 4.7
BIOSSES 24.5 25.2 19.2 13.9
Average 58.2 33.1 16.4 16.1
Table 3: Average word count using GPT-3.5 Turbo for paraphrasing, summarising and extracting keywords.
Dataset Paraphrases Summaries
STSB 0.39 0.52
STS12 0.39 0.49
STS13 0.33 0.41
STS14 0.37 0.47
STS15 0.42 0.53
STS16 0.37 0.40
STS17 0.41 0.59
STS22 0.39 0.21
SICK-R 0.46 0.67
BIOSSES 0.46 0.52
Average 0.40 0.48
Table 4: Average Jaccard Similarity between original texts and paraphrases and summaries generated with GPT-3.5 Turbo respectively.
Augmentation
Original text 1 vs
Augmented text 1
Original text 2 vs
Augmented text 2
Augmented text 1 vs
Augmented text 2
Original text 1 vs
Original text 2
Paraphrasing 85.44 85.24 57.31
$\uparrow$
59.81
$\downarrow$
Summarising 85.53 85.36 57.60
Keyword Extraction 84.27 84.20 59.81
Table 5: Average cosine similarity between original text pairs and augmented texts based on all-mpnet-base-v2 embeddings in %. (The last column shows cosine similarity between the original, non-augmented text pairs.)
Representation learning has emerged as a fundamental technique in natural language processing (NLP). However, the quality and robustness of the embeddings are highly dependent on the richness and diversity of the training data. Recent advancements of generative Large Language Models (LLMs) led to remarkable capabilities in generating human-like text. LLMs were also used for data augmentation by Dai et al. (2023), who applied paraphrasing techniques to train a BERT model (Devlin et al., 2019). Wahle et al. (2022) showed that LLM-generated paraphrases are harder to detect for humans than paraphrases generated with simple techniques like synonym replacement. The integration of generative models with representation learning has been explored by augmenting generative models, e.g., in Retrieval-Augmented Generation (Lewis et al., 2020; Guu et al., 2020).
In contrast, we introduce an approach to augment embedding models that applies generative models using test-time compute: Generatively Augmented Sentence Encoding (GASE). Instead of generating synthetic training data, GASE creates textual variants of an input text through paraphrasing, summarising, or extracting keywords at inference time. A joint embedding is derived by pooling the embeddings of the original text and the generated transformation(s). The underlying hypothesis of our work is that adding textual diversity using generative models increases the ability of current embedding models to model semantics and hence benefits the performance of downstream STS tasks.
Also aiming to propose a training-free method, Lei et al. (2024) introduced Meta-Task Prompting, which employs use-case-specific prompts to generate multiple embeddings via a generative model, after which the embeddings are pooled. For retrieval systems, Gao et al. (2023) and Wang et al. (2023) described query expansion approaches using generative models that generate synthetic documents as a response to a retrieval request. Extending this work, GASE combines generative and embedding models without any assumptions on the type of models (unlike Meta-Task Prompting), nor is it limited to augmentation of retrieval queries such as Gao et al. (2023) and Wang et al. (2023). More recently, GenEOL (Thirukovalluru and Dhingra, 2025), a related approach, was introduced (see Table 11 in the Appendix for a comparison with GASE).
2 Method
GASE performs the following steps (see Figure 1):
Generative Augmentation. We apply $k$ different generative models [^1] Using a single generative model for $k=2$ did not yield textually diverse and meaningful variations. to produce $k$ variations of the original input sequence using exactly one of the following transformations: Paraphrasing, which provides a semantically equivalent but lexically or syntactically different variation of the input text. Summarising, which produces a shorter output text that captures the most important information of the input text. Extracting Keywords, which lists the most relevant words from a given text. Figure 2 provides an example for each. We evaluated $k\in\{0,1,2,3\}$ using GPT-3.5 Turbo (OpenAI, 2024b), Reka-Flash (Reka, 2024), and GPT-4o mini  (OpenAI, 2024a) [^2] For model versions and hyperparameters see Appendix F..
2. Sentence Embedding. Generating embeddings for the original and $k$ synthetic texts with one of $12$ encoders. [^3] See Appendix D for a list incl. references.
3. Pooling. The embeddings of the original text and the $k$ synthetic texts are pooled by computing their arithmetic mean. [^4] Mean pooling dominates max pooling (see Table 12).
We evaluate our approach on the English subtasks of the MTEB STS task [^5] See Appendix E for details on the datasets. using cosine similarity and Spearman’s rank correlation [^6] See  Reimers et al. (2016) for why Spearman’s rank correlation is preferable over Pearson correlation for STS tasks..
3 Results
Embedding model No Aug- gpt-3.5-turbo reka-flash gpt-4o-mini gpt-3.5-turbo gpt-3.5-turbo
mentation + reka-flash + reka-flash
+ gpt-4o-mini
glove.840B.300d 57.86 59.98 60.73 60.04 61.97 62.18
bert-large-cased 60.78 66.29 65.76 65.01 68.01 68.41
all-MiniLM-L6-v2 78.91 80.32 79.74 79.81 80.64 80.76
all-mpnet-base-v2 80.28 81.39 81.01 81.07 81.72 81.85
embed-english-light-v3.0 78.75 80.30 79.72 79.90 80.54 80.66
embed-english-v3.0 81.25 82.10 82.00 81.82 82.53 82.08
voyage-2 82.50 83.38 83.37 83.07 83.80 83.72
voyage-lite-02-instruct 85.95 85.99 85.92 85.97 86.05 86.01
voyage-large-2 83.60 83.92 83.98 83.95 84.12 84.17
voyage-large-2-instruct 84.61 84.80 84.68 84.78 84.86 84.91
mxbai-embed-large-v1 84.63 85.01 84.80 84.90 85.17 85.16
llama-3-8b 71.36 75.28 77.02 74.71 77.85 77.52
Table 6: Average STS scores without and with paraphrase augmentation using different generative models in % (bold: highest score per emb. model, underlined: highest score for $k<=1$ per emb. model).
Table 1 shows the results for different augmentation strategies vs. the respective baseline (see Table 13 in the Appendix for the results per dataset). All embedding models improved their avg. scores through augmentation with the greatest improvements for Llama-3 (+6.496.49pp), BERT (+6.116.11pp), and GloVe (+2.962.96pp). Paraphrase augmentation worked best for most models; however, GloVe improved most with keyword extraction, while BERT, MPNET, and Llama-3 excelled with summarisation.
To validate our approach we compared the augmentation with Keyword Extraction using LLMs to randomly extracting keywords [^7] See Appendix B for implementation details. which demonstrated that LLMs extract semantically more meaningful keywords (Table 2). [^8] Additionally, we compared LLM-based keyword extraction with stopword removal showing that it generates semantically richer extractions (see Table 14 in the Appendix).
Summarising and extracting keywords significantly reduced the average word count across datasets (see Table 3). However, GPT-3.5 Turbo could not maintain STS22’s text length when paraphrasing and at the same time only effectively summarised STS22. Comparing the Jaccard Similarity between original texts and paraphrases on the one hand and original texts and summaries on the other hand, Table 4 shows that paraphrases deviate more from the original texts on all datasets except STS22. At the semantic level (quantified by cosine similarity between embeddings) all augmentation strategies preserve meaning: similarities between the transformed and original texts range from $84.20\%$ to $85.36\%$ (Table 5).
Evaluating the different generative models for paraphrase augmentation we find that using a single generative model GPT-3.5 Turbo performed best for all embedding models, except GloVe, Voyage-Large-2, and Llama-3 (Table 6). [^9] For results per dataset see Table 15 and Table 16 in the Appendix).
Using weighted averages between paraphrases and originals peaked at 25%/75% (excl. BERT) and was worst with paraphrases only (Figure 3).
Applying more than one generative model for augmentation yielded higher scores than a single model (see Table 6) with BERT (+7.637.63pp), Llama-3 (+6.496.49pp) and GloVe (+4.324.32pp) showing the largest gains. Moreover, the overall performance of embedding models with lower non-augmented performance was associated with larger performance improvements through generative augmentation.
Embedding model Augment. STS17 STS22
bert-base-multi-
lingual-cased
None 37.16 29.50
Paraphrase 43.93 30.36
Summarise 47.13 42.03
paraphrase-multiling-
ual-mpnet-base-v2
None 82.51 61.56
Paraphrase 82.95 62.72
Summarise 82.88 66.70
multilingual-
e5-large-instruct
None 83.39 69.64
Paraphrase 83.95 66.48
Summarise 83.60 70.26
voyage-multi-
lingual-2
None 85.14 69.88
Paraphrase 85.73 70.21
Summarise 85.41 71.92
Table 7: Average STS scores for multilingual datasets with and without augmentation with GPT-4o mini in % (bold: highest scores per embedding model and dataset).
Generative Model Size (Qwen2.5)
None 0.5b 1.5b 3b 7b 14b
80.28 79.20 79.79 80.26 80.33 80.79
Table 8: Average STS scores of all-mpnet-base-v2 without and with paraphrase augmentation using Qwen2.5 models with different number of parameters in %.
Embedding
model
Augmen-
tation
PC IR
glove.840B.300d
None 58.41 11.83
Paraphrase 64.64 15.11 (14.12)
bert-large-cased
None 55.07 4.62
Paraphrase 62.36 6.66 (5.62)
all-MiniLM-L6-v2
None 82.34 31.57
Paraphrase 83.80 33.09 (31.82)
all-mpnet-base-v2
None 83.00 33.29
Paraphrase 84.61 35.21 (34.39)
embed-english-
light-v3.0
None 80.18 21.56
Paraphrase 82.88 24.02 (21.63)
embed-english-
v3.0
None 83.54 35.55
Paraphrase 85.12 35.10 (34.34)
voyage-2
None 83.81 29.65
Paraphrase 85.24 30.34 (28.96)
voyage-lite-02-
instruct
None 93.08 35.95
Paraphrase 92.22 36.38 (35.50)
voyage-large-2
None 92.96 33.20
Paraphrase 92.36 32.83 (31.43)
voyage-large-2-
instruct
None 85.21 37.87
Paraphrase 86.22 37.65 (37.44)
mxbai-embed-
large-v1
None 87.16 37.73
Paraphrase 87.38 38.89 (37.94)
llama-3-8b
None 70.10 25.39
Paraphrase 72.78 28.24 (26.39)
Table 9: Avg. Pair Classification (PC) and Information Retrieval (IR) scores with and without paraphrase augmentation (bold: highest scores per embedding model and task). The bracketed IR score shows augmentation applied only to the queries, not the corpus.
Augmentation all-mpnet-base-v2
None 80.28
paraphrases + summaries
82.26
paraphrases + keywords
82.06
summaries + keywords
82.03
paraphrases + summa-
ries + keywords
82.79
Table 10: Average STS scores for different combinations of augmentations generated with GPT-3.5-Turbo for all-mpnet-base-v2 embeddings in % (bold: highest score).
3.1 Extensions
To evaluate the generalisability of GASE we extended our work as follows:
Multilingual. Consistent with prior findings, GASE evaluated on the $2$ multilingual MTEB STS datasets containing $12$ different languages using $3$ multilingual embedding models yields greater improvements for weaker encoders, while summarisation remains most effective for BERT and STS22 (Table 7). [^10] See Table 17 and Table 18 for results per dataset.
Generative model ablation. Ablating the generative model size of Qwen 2.5  (Yang et al., 2025) shows that GASE’s performance correlates with generative model size (Table 8). [^11] See Appendix F for implementation details.
Additional task. We evaluated GASE on Pair Classification (PC) and Information Retrieval (IR). PC was measured with average precision on Twitter SemEval 2015, Twitter URL Corpus, and Sprint Duplicate Questions, while IR used NDCG@10 on NFCorpus. Consistent with STS, GASE improved scores for most models, with gains negatively correlated with baseline performance (Table 9). [^12] See Table 19 in the Appendix for scores per dataset and Muennighoff et al. (2023) for details on score calculations.
Cross-augmentation. Rather than applying a single augmentation strategy across $k$ different generative models, we used a single generative model with multiple augmentation strategies. Tested with all-mpnet-base-v2, using $3$ augmentations the approach outperformed all other (Table 10).
4 Discussion
We believe that embedding models with lower baseline performance benefited more from the semantic diversity induced by augmentation due to their inherent lower capability to model diverse semantics. Similarly, they gained more from the ensemble effect using $k\geq 2$ generative models. Our extension to $11$ non-English languages and the PC and IR tasks did confirm the effectiveness of GASE and also showed a negative association between baseline model performance and improvements through augmentation.
Figure 3: Avg. STS scores for differently weighted averages between embeddings of paraphrases generated with GPT-3.5-Turbo and original texts ($k=1$).
The effectiveness of keyword extraction for GloVe, BERT, and Llama-3 may be explained by the model’s limited capability to handle complex semantics. Therefore, the reduction to keywords might help the model to reduce noise.
Longer texts likely caused summary augmentation to outperform paraphrase augmentation on STS22, as summarising significantly reduced the length. On other datasets, however, summaries were only slightly shorter, diminishing its augmentation value. Moreover, the lower Jaccard Similarity between paraphrases and the original texts indicates greater variation, which in turn makes embeddings more effective. Consistent with this, summaries outperform paraphrases on STS22 (Table 1) - the only dataset where summaries exhibit a lower Jaccard Similarity than paraphrases (Table 4). Our results show pooling original and synthetic embeddings outperforms individual use (Figure 3). Across datasets, tasks, and languages, GASE particularly benefits embedding models with lower baseline performance and shorter texts, while summary augmentation is more effective for longer texts (Table 1). Moreover, GASE requires a certain size of the generative model to be effective (Table 8). For IR, we found that augmenting both the queries and the corpus is necessary (Table 9) which is particularly costly due to document length.
Since GenEOL used different, mostly weaker, embedding models, our results cannot be directly compared with their results.
5 Conclusion
We propose GASE, which augments sentence encoders at inference with generative models for paraphrasing, summarising, or keyword extraction. On MTEB STS, GASE significantly improves lower-performing embedding models and incrementally enhances SOTA models. GASE applies broadly, needing only embedding and generative model outputs. Extensions to $11$ non-English languages and Pair Classification and Information Retrieval as additional tasks, demonstrated consistent findings.
6 Limitations
This study presents several limitations:
  1. 1.
    Our approach significantly increases the computational cost for sentence encoding since each generative augmentation requires an additional inference by the embedding model plus inference using a generative LLM. For a basic estimate of the runtime increase based on STS17 see Table 20. The results highlight GASE’s particular suitability for smaller embedding models, where it yields significant performance gains with minimal encoder runtime.
  2. 2.
    While deterministic embedding models such as GloVe yield consistent results, other models like GPT-3.5 Turbo examined in this research exhibit stochastic behaviour. Consequently, additional experiments are necessary to corroborate our findings, as LLMs can produce diverse outcomes, thereby limiting the conclusiveness of a single experimental run (Reimers and Gurevych, 2018). This caveat is particularly pertinent to observations based on marginal performance differentials.
  3. 3.
    Our experiments demonstrate that the effectiveness of GASE is limited to embedding models with lower baseline performance while encoders with higher baseline performance exhibit minimal to no improvement. Further investigation is needed to understand the underlying reasons for this observation.
  4. 4.
    Our investigation was confined to evaluating GPT-3.5 Turbo, Reka-Flash, GPT-4o mini, and to some limited extent Qwen2.5 for generative augmentation. Alternative LLMs, including the Claude and Gemini model families, may lead to different results.
  5. 5.
    Furthermore, our experimental design was restricted to $k\in\{0,1,2,3\}$. While values of $k>3$ were not explored, we posit that such configurations may be impractical for real-world applications due to prohibitive computational costs.
  6. 6.
    Our experiments for generative model ablation are limited to models with a parameter count $\in\{0.5b,1.5b,3b,7b,14b\}$. Extending the analysis to larger generative models would yield insights into whether the observed pattern holds for models with more than 1414b parameters.
  7. 7.
    Summaries typically comprise $50\%$ or fewer words relative to its source text (Radev et al., 2002). However, due to the concise nature of the texts within the examined STS datasets the summaries generated in this work approximately maintain the original word count. (Except for STS22 which contains longer text sequences.) Hence, the efficacy of augmentation through summarisation may be more pronounced when applied to datasets comprising longer textual inputs (e.g., paragraphs) compared to the predominantly short text sequences examined within the scope of this study.
Acknowledgments
Code development for this work has been assisted by GPT-3.5 Turbo, GPT-4 Omni, Claude 3.5 Sonnet, Gemini Pro 1.5, and Github Copilot.
This research was partially supported by the Horizon Europe project GenDAI (Grant Agreement ID: 101182801) and by the ADAPT Research Centre at Munster Technological University. ADAPT is funded by Taighde Éireann – Research Ireland through the Research Centres Programme and co-funded under the European Regional Development Fund (ERDF) via Grant 13/RC/2106_P2.
We would also like to thank the anonymous reviewers for their valuable feedback and constructive suggestions, which have helped to improve the quality and clarity of this work.
Appendix A Comparison of GASE and GenEOL
A comparison of our approach, GASE, and GenEOL can be found in Table 11.
GASE (ours) GenEOL
Transformations Paraphrasing
Summarising
Extracting keywords
Paraphrasing
Consice paraphrasing
Changing the sentence structure
Entailment
Each one followed by summarisation.
Number of transformations $k\in\{0,1,2,3\}$ $m\in\{0,2,4,8,16,24,32\}$
Pooling Mean pooling over variations generated using $k$ different generative LLMs Mean pooling over variations generated using $m$ different transformations
Embeding models English-language models:
glove.840B.300d
bert-large-cased
all-MiniLM-L6-v2
all-mpnet-base-v2 embed-english-light-v3.0
embed-english-v3.0
voyage-2
voyage-lite-02-instruct
voyage-large-2 voyage-large-2-instruct
mxbai-embed-large-v1
llama-3-8b
Multilingual models:
paraphrase-multilingual-
mpnet-base-v2
multilingual-e5-large-
instruct
voyage-multilingual-2
mistral0.1-7B
llama-2-7B
llama-3-8B
bert-large
Generative models gpt-3.5-turbo-0125
reka-flash-20240226
gpt-4o-mini-2024-07-18
qwen2.5-0.5b
qwen2.5-1.5b
qwen2.5-3b
qwen2.5-7b
qwen2.5-14b
qwen2.5-32b
gpt-3.5-turbo-0125
mistral0.1-I-7B
STS datasets Complete English MTEB for STS:
STSB
STS12
STS13
STS14
STS15
STS16
STS17
STS22
SICK-R
BIOSSES
Multilingual datasets:
STS17
STS22
Subset of English MTEB for STS:
STSB
STS12
STS13
STS14
STS15
STS16
SICK-R
Other tasks Text Pair Classification
Retrieval
Text Pair Classification
Classification
Clustering
Reranking
Table 11: Comparison of GASE and GenEOL.
Based on $7$ STS datasets, Thirukovalluru and Dhingra (2025) report a Spearman’s rank correlation for the weakest baseline model of $47.06$ and $80.37$ for their best model using $32$ generative augmentations. Using the same $7$ datasets, we find that without any augmentation their best performing model is outperformed by $7$ of the $12$ embedding models evaluated in this work (e.g. by all-mpnet-base-v2; see Table 15 and Table 16).
Appendix B Random keyword extraction
Random keyword extraction randomly selects $16.1/58.2\approx 0.28$ of the words from a text sequence. This share is equal to the average share of the number of LLM-extracted keywords and the total word count of a sentence (see Table 3). As this results in a low number of keywords for short sentences, the minimal number of randomly selected keywords was set to $3$. All punctuation has been removed before extracting the keywords.
Appendix C Prompts
Paraphrasing with GPT-3.5 Turbo and GPT-4o mini :
"Rephrase the following text while maintaining its original meaning. If the text contains only a single word, provide a definition or a synomym. When done, check and make sure that the length of the original is approximately maintained. Text:"
Paraphrasing with Reka-Flash:
"Rephrase the following text while maintaining its original meaning. Do not provide multiple alternatives. Before you reply, remove any explanations. Do only reply with the paraphrased text. If the text contains only a single word, provide a definition or a synomym. Text:"
Summarising with GPT-3.5 Turbo:
"summarise the following text. Do not include any meta text and only output the summary. If the text is too short to summarise, paraphrase it instead. Text:"
Extracting keywords with GPT-3.5 Turbo:
"Extract the keywords from the following sentence. Do NOT inlcude any commas or fullstops in your response and do not start your answer with "keywords". Text: "
Model-specific post-processing is performed to remove meta-text not used for augmentation.
Appendix D Embeddings models
Embedding models for the English language:
  • glove.840B.300d (Pennington et al., 2014)
  • bert-large-cased (Devlin et al., 2019)
  • all-MiniLM-L6-v2 Sentence Transformers (2024)
  • all-mpnet-base-v2 Sentence Transformers (2024)
  • embed-english-light-v3.0 Cohere (2024)
  • embed-english-v3.0 Cohere (2024)
  • voyage-2 (Voyage AI, 2024)
  • voyage-lite-02-instruct (Voyage AI, 2024)
  • voyage-large-2 (Voyage AI, 2024)
  • voyage-large-2-instruct (Voyage AI, 2024)
  • mxbai-embed-large-v1 Mixedbread (2024)
  • llama-3-8b Meta (2024)
For llama-3-8b we used the following KEEOL Zhang et al. (2024) prompt:
The essence of a sentence is often captured by its main subjects and actions, while descriptive terms provide additional but less central details. With this in mind, this sentence: “input_text” means in one word:
Following Zhang et al. (2024), we used the penultimate layer of llama-3-8b to extract the embeddings.
Multilingual embedding models:
  • paraphrase-multilingual-mpnet-
    base-v2
     Reimers and Gurevych (2020)
  • intfloat/multilingual-e5-
    large-instruct
     Wang et al. (2024)
  • voyage-multilingual-2 Voyage AI (2024)
Appendix E Datasets
The following datasets from the MTEB STS leaderboard (Hugging Face, 2024) were used:
  • STSB (Cer et al., 2017)
  • STS12 (Agirre et al., 2012)
  • STS13 (Agirre et al., 2013)
  • STS14 (Agirre et al., 2014)
  • STS15 (Agirre et al., 2015)
  • STS16 (Agirre et al., 2016)
  • STS17 (Cer et al., 2017)
  • STS22 (Chen et al., 2022)
  • SICK-R Marelli et al. (2014)
  • BIOSSES (Soğancıoğlu et al., 2017)
The original MTEB paper (Muennighoff et al., 2023) also included the STS11 dataset in Figure 1 of their paper where the authors specify the MTEB datasets, but excluded it in the evaluations. Similarly, the MTEB Leaderboard (Hugging Face, 2024) does not include STS11 either. To make our results comparable to other models evaluated on MTEB we therefore excluded STS11 from our evaluations. All data has been obtained from the MTEB repository on the Hugging Face website. [^13] https://huggingface.co/mteb/datasets
The following datasets were used for our extensions:
  • Twitter SemEval 2015 (Xu et al., 2015)
  • Twitter URL-Corpus (Lan et al., 2017)
  • Sprint duplicate questions (Shah et al., 2018)
  • NFCorpus (Boteva et al., 2016)
Appendix F Generative Model Versions, Hyperparameters, and Hardware Configuration
The following versions have been used for generative augmentation:
  • GPT-3.5 Turbo: gpt-3.5-turbo-0125
  • Reka-Flash: reka-flash-20240226
  • GPT-4o mini: gpt-4o-mini-2024-07-18
  • Qwen2.5 (all models running via Ollama with quantization Q4_K_M):
    • qwen2.5:0.5b
    • qwen2.5:1.5b
    • qwen2.5:3b
    • qwen2.5:7b
    • qwen2.5:14b
GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini have been accessed through the respective APIs with temperature set to 0 and top_p to $1$ to make the experiments as reproducible as possible. Where applicable, a random seed of $1337$ was used. Other hyperparameters were used with default values.
All Qwen2.5 model runs were conducted with Ollama on a local computer with Q4_K_M quantisation and the following specification:
  • GeForce RTX 3090 MSI Gaming X Trio 24GB GPU
  • AMD Ryzen 9 9950X CPU
  • DDR5-6400 64GB RAM
Appendix G Additional Experimental Results
Embedding model Pooling STS Score
glove.840B.300d
mean 59.98
max 59.02
bert-large-cased
mean 66.29
max 65.16
all-MiniLM-L6-v2
mean 80.32
max 79.88
all-mpnet-base-v2
mean 81.39
max 81.03
embed-english-
light-v3.0
mean 80.30
max 79.74
embed-english-v3.0
mean 82.10
max 81.62
voyage-2
mean 83.38
max 82.69
voyage-lite-02-
instruct
mean 85.99
max 85.53
voyage-large-2
mean 83.92
max 83.48
voyage-large-2-
instruct
mean 84.80
max 84.41
mxbai-embed-
large-v1
mean 85.01
max 84.69
llama-3-8b
mean 75.28
max 75.07
Table 12: Average STS scores across tasks for different pooling methods in % (bold: best per embedding model).
Embedding model Augmentation Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
glove.840B.300d none 57.86 50.73 57.50 70.98 60.69 70.85 63.85 62.05 54.08 55.42 32.45
paraphrasing 59.98 56.78 57.33 71.17 61.16 71.52 67.66 62.28 57.78 58.92 35.21
summarising 60.64 57.67 55.00 70.24 61.18 71.64 66.81 67.08 61.21 60.92 34.65
extracting keywords 60.82 59.26 58.09 71.12 60.34 72.44 65.31 75.92 56.94 55.34 33.50
bert-large-cased none 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
paraphrasing 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.79 64.56
summarising 66.89 68.23 51.39 70.58 61.98 74.27 70.60 74.48 63.47 70.74 63.23
extracting keywords 62.62 63.03 53.50 63.32 58.51 72.62 68.13 71.99 54.64 58.24 62.22
all-MiniLM-L6-v2 none 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
paraphrasing 80.32 82.33 73.33 83.33 77.85 86.23 79.88 87.02 68.47 79.63 85.07
summarising 79.90 82.17 72.14 82.89 77.24 86.34 78.99 87.70 69.27 79.86 82.41
extracting keywords 77.96 81.99 73.29 81.14 75.77 85.87 78.70 87.45 64.57 70.92 79.88
all-mpnet-base-v2 none 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
paraphrasing 81.39 84.10 74.04 85.52 80.52 86.82 81.15 89.08 69.01 81.25 82.39
summarising 81.48 83.94 73.76 85.54 79.78 86.73 80.20 90.41 70.12 81.32 82.96
extracting keywords 79.95 82.56 74.82 84.82 78.74 86.31 79.01 90.23 66.19 77.07 79.74
embed-english-light-v3.0 none 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
paraphrasing 80.30 84.65 74.65 80.73 79.41 85.45 81.14 87.45 68.42 78.75 82.36
summarising 80.21 84.08 73.45 81.36 78.41 86.26 81.52 88.19 68.97 79.08 80.79
extracting keywords 78.53 82.81 75.03 80.17 78.29 84.14 76.38 88.14 67.27 74.45 78.64
embed-english-v3.0 none 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
paraphrasing 82.10 86.69 76.37 83.71 81.29 87.96 83.95 88.47 68.63 78.21 85.75
summarising 82.08 86.47 74.40 84.41 80.76 88.49 83.55 89.08 70.29 78.06 85.25
extracting keywords 81.26 85.50 74.77 84.66 79.98 87.72 81.09 88.75 66.91 77.53 85.68
voyage-2 none 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
paraphrasing 83.38 87.17 79.23 87.13 81.67 88.57 86.27 89.07 66.80 80.33 87.56
summarising 82.74 86.51 76.81 86.25 80.89 88.46 84.88 89.10 68.50 80.48 85.55
extracting keywords 81.21 85.71 77.76 86.17 79.63 87.49 84.37 88.66 65.29 72.06 85.03
voyage-lite-02-instruct none 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
paraphrasing 85.99 88.81 84.47 89.05 86.68 89.52 86.53 86.79 79.08 79.09 89.89
summarising 85.43 87.93 82.77 88.06 85.05 88.98 85.49 86.70 79.87 79.16 90.32
extracting keywords 83.91 86.75 83.82 87.97 84.47 88.98 84.55 86.28 76.60 70.77 88.88
voyage-large-2 none 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
paraphrasing 83.92 87.62 79.18 87.66 83.60 88.91 86.00 90.22 64.69 81.00 90.29
summarising 83.66 87.56 77.83 86.93 82.88 88.93 84.88 90.49 65.90 81.09 90.15
extracting keywords 81.98 85.92 78.00 86.45 81.33 88.28 84.44 89.90 63.68 72.50 89.33
voyage-large-2-instruct none 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
paraphrasing 84.80 89.31 77.61 88.40 86.62 91.17 86.20 89.58 67.72 82.83 88.58
summarising 84.54 88.84 76.27 88.29 85.56 91.06 85.24 89.40 68.66 83.28 88.78
extracting keywords 83.40 87.38 76.63 88.58 85.55 90.51 84.03 88.57 65.73 76.58 90.46
mxbai-embed-large-v1 none 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
paraphrasing 85.01 89.09 80.83 89.43 85.88 89.46 86.94 88.75 70.03 82.67 87.02
summarising 84.45 88.60 78.51 89.07 84.82 89.09 85.96 88.43 70.36 82.82 86.82
extracting keywords 83.42 87.43 78.64 88.75 83.82 88.44 84.84 87.40 67.83 78.17 88.90
llama-3-8b none 71.36 79.79 62.81 79.08 69.72 78.49 80.88 84.12 33.39 76.27 69.04
paraphrasing 75.28 81.37 67.03 81.29 72.88 80.91 82.60 83.04 50.55 78.25 74.84
summarising 76.74 80.39 65.02 81.18 74.51 82.07 81.00 84.48 59.44 79.36 79.92
extracting keywords 74.95 78.57 67.39 82.68 74.74 81.67 80.31 85.71 51.79 71.05 75.60
Table 13: Scores per dataset with and without generative augmentation using GPT-3.5 Turbo in % (bold: highest scores).
Embedding
model
Augment-
tation
Stop-
word
removal
Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
glove.840B.
300d
none no 57.86 50.73 57.50 70.98 60.69 70.85 63.85 62.05 54.08 55.42 32.45
yes 60.07 57.13 57.43 70.91 60.68 70.86 63.84 75.83 54.66 54.97 34.39
para-
phrasing
no 59.98 56.78 57.33 71.17 61.16 71.52 67.66 62.28 57.78 58.92 35.21
yes 65.43 67.05 59.98 72.72 63.75 73.70 69.98 75.01 61.01 61.94 49.17
bert-large-cased none no 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
yes 53.08 50.11 41.37 53.10 50.11 63.68 59.66 63.73 49.69 48.75 50.54
para-
phrasing
no 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.79 64.56
yes 60.25 61.42 50.45 58.43 55.20 69.75 67.19 67.98 57.01 58.25 56.82
all-MiniLM-
L6-v2
none no 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
yes 75.36 78.80 68.29 77.62 73.05 84.35 74.92 87.05 65.74 65.58 78.24
para-
phrasing
no 80.32 82.33 73.33 83.33 77.85 86.23 79.88 87.02 68.47 79.63 85.07
yes 77.78 80.69 71.61 81.15 75.92 84.86 77.51 85.89 66.36 70.61 83.21
all-mpnet-base-
v2
none no 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
yes 75.51 78.32 65.71 79.49 73.99 83.98 73.28 88.98 65.82 67.01 78.47
para-
phrasing
no 81.39 84.10 74.04 85.52 80.52 86.82 81.15 89.08 69.01 81.25 82.39
yes 78.20 80.87 70.76 83.09 77.53 85.00 76.72 87.05 66.28 71.99 82.70
embed-english-
light-v3.0
none no 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
yes 74.36 78.16 68.36 76.23 73.84 81.67 71.69 88.18 66.87 64.32 74.34
para-
phrasing
no 80.30 84.65 74.65 80.73 79.41 85.45 81.14 87.45 68.42 78.75 82.36
yes 77.58 81.43 73.12 79.15 76.81 83.77 76.89 87.55 67.81 70.40 78.87
embed-english-
v3.0
none no 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
yes 76.70 79.52 67.15 80.84 75.37 83.76 74.40 88.51 66.99 66.56 83.95
para-
phrasing
no 82.10 86.69 76.37 83.71 81.29 87.96 83.95 88.47 68.63 78.21 85.75
yes 79.26 83.00 72.79 83.09 78.72 85.21 79.03 87.36 68.03 71.45 83.96
voyage-2 none no 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
yes 78.11 81.52 71.71 82.73 77.23 85.59 79.12 87.91 66.40 67.22 81.64
para-
phrasing
no 83.38 87.17 79.23 87.13 81.67 88.57 86.27 89.07 66.80 80.33 87.56
yes 80.42 84.16 76.32 84.89 79.29 86.33 82.93 87.58 67.55 71.42 83.69
voyage-lite-02-
instruct
none no 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
yes 81.35 83.24 79.33 86.13 82.37 87.95 80.01 85.84 74.72 66.73 87.14
para-
phrasing
no 85.99 88.81 84.47 89.05 86.68 89.52 86.53 86.79 79.08 79.09 89.89
yes 82.88 85.57 81.64 87.14 83.29 87.58 83.46 85.18 76.02 70.83 88.11
voyage-large-2 none no 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
yes 79.39 82.25 72.85 82.94 78.99 86.54 79.68 89.63 65.38 67.42 88.25
para-
phrasing
no 83.92 87.62 79.18 87.66 83.60 88.91 86.00 90.22 64.69 81.00 90.29
yes 81.19 84.57 76.72 84.73 80.84 87.17 82.95 88.78 65.43 71.64 89.04
voyage-large-2-
instruct
none no 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
yes 81.35 83.24 79.33 86.13 82.37 87.95 80.01 85.84 74.72 66.73 87.14
para-
phrasing
no 84.80 89.31 77.61 88.40 86.62 91.17 86.20 89.58 67.72 82.83 88.58
yes 81.88 85.70 74.85 87.76 83.91 88.69 82.40 88.41 67.66 71.72 87.73
mxbai-embed-
large-v1
none no 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
yes 79.54 83.54 72.88 85.73 81.33 86.72 78.69 86.36 67.41 68.76 83.99
para-
phrasing
no 85.01 89.09 80.83 89.43 85.88 89.46 86.94 88.75 70.03 82.67 87.02
yes 81.36 86.15 78.11 87.20 83.03 86.93 78.69 86.40 68.47 72.70 85.97
llama-3-8b
none no 71.36 79.79 62.81 79.08 69.72 78.49 80.88 84.12 33.39 76.27 69.04
yes 69.50 73.48 59.66 77.37 69.66 77.17 72.94 83.00 50.94 62.15 68.58
para-
phrasing
no 75.28 81.37 67.03 81.29 72.88 80.91 82.60 83.04 50.55 78.25 74.84
yes 74.57 78.29 66.98 80.90 73.61 79.63 78.76 83.86 61.88 68.92 72.90
Table 14: Scores per dataset with and without stopword removal using either no generative augmentation or paraphrase augmentation with GPT-3.5 Turbo in %.
Embedding model Gen model Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
glove.840B.300d none 57.86 50.73 57.50 70.98 60.69 70.85 63.85 62.05 54.08 55.42 32.45
gpt-3.5-turbo 59.98 56.78 57.33 71.17 61.16 71.52 67.66 62.28 57.78 58.92 35.21
reka-flash 60.73 55.02 59.53 71.09 61.22 72.81 65.20 64.27 60.11 61.65 36.43
gpt-4o-mini 60.04 55.09 60.12 72.65 62.45 71.60 66.49 63.39 55.48 60.36 32.76
gpt-3.5-turbo
+ reka-flash
61.97 59.46 60.06 72.01 62.68 72.46 67.62 65.43 61.46 61.51 36.95
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
62.18 59.99 61.79 72.88 64.04 72.03 67.38 64.75 60.76 61.79 36.42
bert-large-cased none 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
gpt-3.5-turbo 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.80 64.56
reka-flash 65.76 65.20 52.92 68.55 59.77 72.67 71.80 71.46 58.93 70.09 66.19
gpt-4o-mini 65.01 65.37 54.27 68.60 58.18 72.26 71.12 73.54 55.72 70.50 60.54
gpt-3.5-turbo
+ reka-flash
68.01 69.39 56.86 70.04 62.30 74.61 73.86 74.52 60.49 71.32 66.72
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
68.41 69.40 59.01 71.25 62.60 74.13 73.86 74.99 60.35 71.74 66.73
all-MiniLM-L6-v2 none 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
gpt-3.5-turbo 80.32 82.33 73.33 83.33 77.85 86.23 79.88 87.02 68.47 79.63 85.07
reka-flash 79.74 81.30 73.52 82.42 76.83 85.72 79.02 87.01 68.83 79.66 83.11
gpt-4o-mini 79.81 81.97 74.03 83.31 76.70 86.03 79.19 87.69 67.01 80.08 82.11
gpt-3.5-turbo
+ reka-flash
80.64 82.30 74.59 83.72 78.11 86.20 79.95 87.23 69.23 80.12 84.97
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
80.76 82.31 75.16 84.29 78.22 86.18 79.98 87.31 68.84 80.34 85.02
all-mpnet-base-v2 none 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
gpt-3.5-turbo 81.39 84.10 74.04 85.52 80.52 86.82 81.15 89.08 69.01 81.25 82.39
reka-flash 81.01 83.72 74.27 85.12 79.33 86.41 80.41 90.17 68.42 81.37 80.87
gpt-4o-mini 81.07 83.75 74.19 85.19 79.28 86.45 80.47 90.89 67.81 81.23 81.43
gpt-3.5-turbo
+ reka-flash
81.72 84.43 75.46 86.12 80.63 87.00 81.21 89.37 69.33 81.55 82.13
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
81.85 84.42 75.96 86.37 80.72 86.92 81.33 89.68 69.05 81.50 82.51
embed-english-light-v3.0 none 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
gpt-3.5-turbo 80.30 84.65 74.65 80.73 79.41 85.45 81.14 87.45 68.42 78.75 82.36
reka-flash 79.72 83.51 75.74 78.60 78.38 84.63 80.09 87.98 69.06 78.76 80.42
gpt-4o-mini 79.90 84.14 75.76 79.65 77.89 85.12 79.83 87.90 67.91 78.68 82.10
gpt-3.5-turbo
+ reka-flash
80.54 84.48 76.45 80.61 79.65 85.54 81.21 87.60 69.23 79.06 81.55
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
80.66 84.59 77.10 81.00 79.75 85.68 81.16 87.70 68.90 79.09 81.61
embed-english-v3.0 none 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
gpt-3.5-turbo 82.10 86.69 76.37 83.71 81.29 87.96 83.95 88.47 68.63 78.21 85.75
reka-flash 82.00 86.31 76.91 82.22 80.48 87.82 83.69 88.93 69.70 78.06 85.90
gpt-4o-mini 81.82 86.52 76.80 82.57 80.11 87.77 83.72 89.42 68.36 77.90 85.05
gpt-3.5-turbo
+ reka-flash
82.53 86.68 77.88 83.56 81.64 88.05 84.13 88.59 69.87 78.44 86.48
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
82.08 86.69 78.49 83.60 81.65 87.97 84.10 88.76 69.67 78.35 86.36
Table 15: Scores per dataset with and without paraphrase augmentation using GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini in % (bold: highest scores) (part 1).
Embedding model Gen model Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
voyage-2 none 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
gpt-3.5-turbo 83.38 87.17 79.23 87.13 81.67 88.57 86.27 89.07 66.80 80.33 87.56
reka-flash 83.37 86.93 79.65 87.83 81.79 88.48 86.18 88.98 67.42 80.67 85.72
gpt-4o-mini 83.07 86.89 79.59 87.27 81.23 88.46 85.81 89.37 65.25 80.85 86.02
gpt-3.5-turbo
+ reka-flash
83.80 87.27 80.62 87.99 82.40 88.51 86.40 88.77 67.94 80.87 87.28
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
83.72 87.14 81.06 88.14 82.50 88.40 86.14 88.81 66.91 81.00 87.08
voyage-lite-02-instruct none 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
gpt-3.5-turbo 85.99 88.81 84.47 89.05 86.68 89.52 86.53 86.79 79.08 79.09 89.89
reka-flash 85.92 88.24 84.88 88.79 86.58 89.46 86.28 86.74 78.76 78.97 90.53
gpt-4o-mini 85.97 88.25 85.24 89.06 86.62 89.57 85.99 86.90 78.59 79.45 90.08
gpt-3.5-turbo
+ reka-flash
86.05 88.74 84.83 89.14 86.69 89.36 86.57 86.72 78.81 79.34 90.34
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
86.01 88.48 84.95 89.26 86.74 89.25 86.32 86.75 78.31 79.50 90.55
voyage-large-2 none 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
gpt-3.5-turbo 83.92 87.62 79.18 87.66 83.60 88.91 86.00 90.22 64.69 81.00 90.29
reka-flash 83.98 87.32 79.52 88.25 83.49 88.77 85.90 90.40 64.46 81.35 90.37
gpt-4o-mini 83.95 87.54 80.02 87.91 83.22 89.11 85.68 90.47 63.99 81.54 90.08
gpt-3.5-turbo
+ reka-flash
84.12 87.56 80.05 88.31 83.98 88.76 85.97 90.04 64.84 81.55 90.14
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
84.17 87.49 80.53 88.42 84.00 88.73 85.77 89.82 64.69 81.75 90.49
voyage-large-2-instruct none 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
gpt-3.5-turbo 84.80 89.31 77.61 88.40 86.62 91.17 86.20 89.58 67.72 82.83 88.58
reka-flash 84.68 88.81 78.41 88.30 86.40 90.95 85.42 89.71 67.04 83.22 88.52
gpt-4o-mini 84.78 88.83 78.71 88.43 86.45 91.17 85.89 89.14 67.09 83.19 88.87
gpt-3.5-turbo
+ reka-flash
84.86 89.16 79.11 88.56 86.67 90.97 85.86 89.67 67.56 83.00 88.08
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
84.91 88.98 79.79 88.59 86.69 90.89 85.83 89.23 67.62 82.97 88.51
mxbai-embed-large-v1 none 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
gpt-3.5-turbo 85.01 89.09 80.83 89.43 85.88 89.46 86.94 88.75 70.03 82.67 87.02
reka-flash 84.80 88.93 81.04 89.36 85.51 89.27 86.48 89.04 68.67 82.66 87.07
gpt-4o-mini 84.90 88.95 81.08 89.76 85.37 89.43 86.58 89.14 69.10 82.73 86.89
gpt-3.5-turbo
+ reka-flash
85.17 89.16 81.96 89.37 85.96 89.39 86.79 88.97 69.51 82.72 87.81
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
85.16 89.13 82.24 89.54 85.98 89.36 86.70 88.92 69.44 82.65 87.63
llama-3-8b none 71.36 79.79 62.81 79.08 69.72 78.49 80.88 84.12 33.39 76.27 69.04
gpt-3.5-turbo 75.28 81.37 67.03 81.29 72.88 80.91 82.60 83.04 50.55 78.25 74.84
reka-flash 77.02 80.44 68.26 82.29 74.75 81.65 82.53 85.77 58.68 78.43 77.38
gpt-4o-mini 74.71 80.38 68.15 82.09 73.39 81.20 81.87 84.60 41.34 78.82 75.23
gpt-3.5-turbo
+ reka-flash
77.85 81.68 69.82 82.99 75.64 82.10 83.27 84.46 59.90 78.91 79.68
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
77.52 81.41 70.96 83.47 75.76 81.99 82.84 84.14 56.86 78.99 78.77
Table 16: Scores per dataset with and without paraphrase augmentation using GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini in % (bold: highest scores) (part 2).
Dataset paraphrase-multilingual- mpnet-base-v2 multilingual- e5-large-instruct voyage-multilingual-2
None
Paraphrase
Augmentation
None
Paraphrase
Augmentation
None
Paraphrase
Augmentation
STS17_en-en 86.99 86.01 87.64 87.32 90.81 90.20
STS17_ar-ar 79.09 81.04 81.61 83.36 80.53 81.58
STS17_en-ar 80.85 81.76 78.66 79.50 81.59 83.48
STS17_en-de 83.28 83.81 85.09 85.05 88.09 88.11
STS17_en-tr 74.90 75.44 77.04 78.88 74.12 77.09
STS17_es-en 86.11 86.25 84.72 85.35 88.24 87.74
STS17_es-es 85.14 85.00 88.09 87.87 88.71 88.63
STS17_fr-en 81.17 82.47 83.26 84.09 87.10 87.65
STS17_it-en 84.24 84.20 84.94 85.00 88.98 89.43
STS17_ko-ko 83.29 83.83 81.58 82.47 80.17 80.69
STS17_nl-en 82.51 82.65 84.65 84.59 88.22 88.38
Table 17: Multilingual STS17 Scores with and without paraphrase augmentation using GPT-4o mini in % (bold: highest score per embedding model and dataset).
Dataset paraphrase-multilingual- mpnet-base-v2 multilingual-e5- large-instruct voyage-multilingual-2
None
Paraphrase
Augmentation
None
Paraphrase
Augmentation
None
Paraphrase
Augmentation
STS_ar-ar 64.77 56.36 81.61 65.02 59.65 60.19
STS_de-de 62.30 48.16 78.66 62.00 63.01 63.08
STS_de-en 62.85 57.03 85.09 61.72 62.87 62.56
STS_de-fr 64.19 62.68 77.04 65.52 64.88 64.12
STS_de-pl 56.01 41.26 84.72 56.97 54.14 56.35
STS_es-es 73.27 65.03 88.09 73.07 75.32 75.87
STS_es-en 80.45 74.22 83.26 79.99 82.43 83.79
STS_es-it 78.71 56.99 84.94 78.76 79.51 79.78
STS_fr-fr 82.34 76.56 81.58 82.35 81.36 81.35
STS_fr-pl 73.25 84.52 84.65 28.17 73.25 73.25
STS_it-it 80.69 62.82 64.77 80.23 79.92 80.15
STS_pl-pl 46.83 39.36 62.30 45.29 46.30 46.00
STS_pl-en 76.42 76.90 62.85 75.61 80.01 80.81
STS_ru-ru 70.26 63.33 64.19 70.74 68.35 68.91
STS_tr-tr 69.56 63.49 56.01 70.16 67.98 68.34
STS_zh-zh 66.92 66.89 73.27 68.53 72.22 72.49
STS_zh-en 69.56 70.68 80.45 66.06 76.81 76.53
Table 18: Multilingual STS22 Scores with and without paraphrase augmentation using GPT-4o mini in % (bold: highest score per embedding model and dataset).
Embedding model Augmen- tation Text Pair Classification
Twitter Sem-
Eval 2015
Twitter URL-
Corpus
Sprint dupli-
cate questions
glove.840B.300d
none 43.81 63.09 68.33
paraphrase 50.04 70.49 73.39
bert-large-cased
none 47.27 69.81 48.15
paraphrase 57.21 76.28 53.61
all-MiniLM-L6-v2
none 67.86 84.61 94.55
paraphrase 71.50 84.70 95.20
all-mpnet-base-v2
none 73.85 84.99 90.15
paraphrase 77.34 85.08 91.42
embed-english-
light-v3.0
none 68.26 83.75 88.54
paraphrase 72.76 84.49 91.40
embed-english-v3.0
none 73.43 84.98 92.21
paraphrase 76.58 85.45 93.33
voyage-2
none 71.29 86.25 93.88
paraphrase 74.83 86.19 94.70
voyage-lite-02-
instruct
none 91.38 89.14 98.35
paraphrase 90.31 88.37 98.39
voyage-large-2
none 75.10 86.57 93.96
paraphrase 76.82 86.58 95.25
voyage-large-2-
instruct
none 92.87 90.10 96.28
paraphrase 90.78 89.09 96.78
mxbai-embed-
large-v1
none 78.55 86.12 96.82
paraphrase 79.17 86.03 96.95
llama-3-8b
none 69.70 82.73 57.87
paraphrase 73.00 84.26 61.09
Table 19: Performance on Text Pair Classification with and without paraphrase augmentation using GPT-3.5-Turbo in %.
Appendix H Runtime analysis
Augmentation Average runtime [sec] Stand. deviation of runtime [sec]
None 2.30 0.40
paraphrase 132.57 15.87
summarise 140.48 14.15
keywords 127.17 18.16
Table 20: Average runtime and corresponding standard deviation without and with generative augmentation using GPT-3.5-Turbo and, as an encoder, all-mpnet-base-v2 on STS17 (based on 55 runs each, in seconds).
To provide a basic estimate of the additional runtime introduced by GASE over a non-augmented baseline, we performed five runs using GPT-3.5-Turbo for generation and all-mpnet-base-v2 for embeddings and report the mean and standard deviation. Because the difference in runtimes between the augmented and non-augmented setups was substantial and the standard deviation across runs was small, we did not carry out statistical significance testing (see Table 20).
All runtime experiments were conducted using Ollama on a local computer with the following specification: GeForce RTX 3090 MSI Gaming X Trio 24GB GPU, AMD Ryzen 9 9950X CPU, DDR5-6400 64GB RAM.